OpenStudy (anonymous):

The function f(x) = 6x + 5 is defined over the interval [-1, 4]. If the interval is divided into n equal parts, what is the value of the right end point of the first rectangle?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

\[x_{1}=-1+\frac{ 5 }{ n }\]

OpenStudy (anonymous):

\[x _{1}=4+\frac{ 5 }{ n }\]

OpenStudy (anonymous):

\[x _{1}=-4+\frac{ 7 }{ n }\]

OpenStudy (anonymous):

\[x _{1}=3+\frac{ 7 }{ n }\]

OpenStudy (anonymous):

\[x _{1}=3+\frac{ 5 }{ n }\]

OpenStudy (anonymous):

@Michele_Laino